OpenStudy (anonymous):

Evaluate the polynomial 6x – y for x = –3 and y = 2.

OpenStudy (anonymous):

where ever you see y sub in a 2 and where ever you see x sub in -3 so in this case we have 6(-3) - (2) = -18 - 2 = -20

OpenStudy (anonymous):

put the given values in for the variables 6(-3) - (2) 6(-3) means 6 times -3, which is -18 so: -18 -2 = -20

OpenStudy (anonymous):

Ah ok! Thank you!!